• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 20:02:51 on August16,2006

    Entry number 180425

    keyword=shift 1/2 way summary

    5:15 Left spectrometer can be moved!
    5:45 call for beam,
    6:05 spectometer is ready, asked 20 mkA
    run 2823 / 21166 for C-optics 20 mkA
    run 2824 / 21167 for 4 cm dummy 20 mkA
    run 2825 / 21168 for 15 cm dummy 20 mkA
    run 2826 / 21169 for Tl 20 mkA
    tune LH2 target for 20 mkA settings
    run 2827 / 21170 for LH2 15 cm 5% radiator 10 mkA
    run 2828 / 21171 for LH2 15 cm 0% 10 mkA
    run 2829 / 21172 for LH2 15 cm 0% radiator 20 mkA